# Augmenting LLMs for General Time Series Understanding and Prediction

### v1

@@ C.4 Synthetic Data Generation
We created three synthetic variants of the PTB-XL dataset to expand its applicability beyond traditional classification tasks. Using Llama 3.3 70B Instruct [Grattafiori et al., 2024] with controlled generation parameters (temperature=0.7, top-k=40), we systematically generated textual content paired with the existing ECG recordings.
@@
To establish and evaluate basic time series manipulation capabilities, we created a synthetic instruction-following dataset using time series sampled from the GIFT-Eval pretraining split. Employing various state-of-the-art LLMs including Gemini 1.5 Pro [Team et al., 2023], we generated diverse instruction-response pairs across four fundamental tasks that serve as building blocks for more complex time series operations.
@@
To generate high-quality responses, we employed a diverse ensemble of state-of-the-art language models including text-only models (Llama 3.3 70B [Grattafiori et al., 2024]), vision-language models (Qwen2-VL-72B [Wang et al., 2024]), and frontier models (Gemini 2.0 Flash [Team et al., 2023]). While synthetic responses may contain some inaccuracies, recent findings [Bansal et al., 2024] suggest that larger quantities of moderate-quality training data can be more effective than smaller sets of perfect examples for LLM training. The resulting scale and diversity of our dataset provides comprehensive coverage of different analytical scenarios, making it particularly suitable for training models in time series analysis and reasoning tasks.
Appendix D Data Samples
@@ Conclusion:
Appendix E Broader Impacts

### v2

@@ B.4 Synthetic Data Generation
We created three synthetic variants of the PTB-XL dataset to expand its applicability beyond traditional classification tasks. Using Llama 3.3 70B Instruct [27] with controlled generation parameters (temperature=0.7, top-k=40), we systematically generated textual content paired with the existing ECG recordings.
@@
To establish and evaluate basic time series manipulation capabilities, we created a synthetic instruction-following dataset using time series sampled from the GIFT-Eval pretraining split. Employing various state-of-the-art LLMs including Gemini 1.5 Pro [51], we generated diverse instruction-response pairs across four fundamental tasks that serve as building blocks for more complex time series operations.
@@
To generate high-quality responses, we employed a diverse ensemble of state-of-the-art language models including text-only models (Llama 3.3 70B [27]), vision-language models (Qwen2-VL-72B [26]), and frontier models (Gemini 2.0 Flash [51]). While synthetic responses may contain some inaccuracies, recent findings [52] suggest that larger quantities of moderate-quality training data can be more effective than smaller sets of perfect examples for LLM training. The resulting scale and diversity of our dataset provides comprehensive coverage of different analytical scenarios, making it particularly suitable for training models in time series analysis and reasoning tasks.
Appendix C Data Samples
@@ Conclusion:
Appendix D Broader Impacts